OpenStudy (znappydooz):

Can anybody help in this question? The following is an indirect proof of the Addition Property of Equality: For real numbers a, b, and c, if a = b, then a + c = b + c. Assume____________. According to the given information, a = b. By the Subtraction Property of Equality, one can subtract the same number from both sides of an equation without changing the equation. Therefore, a + c − c ≠ b + c − c. Through subtraction, the c's cancel and a ≠ b. This contradicts the given information so a + c = b + c. Which statement accurately completes the proof? a + c = b + c a + c ≠ b + c a

OpenStudy (anonymous):

i can probly help you

OpenStudy (znappydooz):

Lol I would really appreciate it! Thnx!

OpenStudy (znappydooz):

here r the rest of teh answers...

OpenStudy (anonymous):

its a

OpenStudy (znappydooz):

a + c = b + c a + c ≠ b + c a ≠ b b = c

OpenStudy (anonymous):

this one is tough

OpenStudy (znappydooz):

How do u kno it s A?

OpenStudy (anonymous):

oh i already said it was a

OpenStudy (anonymous):

read the question carefully and it will give you the answer

OpenStudy (znappydooz):

See, i was leaning more for B

OpenStudy (anonymous):

oh it looks like you are right

OpenStudy (anonymous):

sorry

OpenStudy (znappydooz):

So you think its B too?

OpenStudy (anonymous):

yes

OpenStudy (znappydooz):

Alright! Thnx, i'll see if it's right...

OpenStudy (znappydooz):

yes, it was correct! Thnx so much!

OpenStudy (anonymous):

is it right

OpenStudy (anonymous):

no prob